# Unlocking Scaling Law in Industrial Recommendation Systems with a Three-step Paradigm based Large User Model

### v1

@@ 4.3. Step 3: Knowledge Utilization in DLRMs
To enhance existing DLRMs, we propose two strategies (see Figure 3 (c)): (1) Direct Feature Incorporation: We integrate the outputs $o^{i}_{q_{n}}$ and the embedding $e^{i}_{i}$ of target item $i$ as fixed additional features into the DLRMs. This approach leverages the rich representations derived from the next-condition items directly. (2) Interest Matching via Similarity Measurement: We assess the alignment between the target item $i$ and user interests by computing the similarity $sim(o^{i}_{q_{n}},e^{i}_{i})$ (as defined in Equ 2). This similarity score quantifies how well the target item matches the context provided by the next-condition items.
@@ 4.4. Discussion
$\bullet$ Addressing Limitations 1: To address the first limitation, we design a generative-to-discriminative process. In Step 1, the joint distribution $p(c_{1},i_{1},c_{2},i_{2},...,c_{L},i_{L})$ is constructed through generative learning. Subsequently, the design of ’next-condition-item prediction’ enables the triggering of relevant knowledge for RecSys tasks. Finally, discriminative learning in Step 3 is performed to meet the requirements of discriminative applications.
$\bullet$ Addressing Limitations 2: The model training and serving in Step 3 naturally support the efficiency requirements in industrial applications, as the backbone of the model follows the DLRM-style architecture. Additionally, the computational costs associated with Steps 1 and 2 are minimal due to the decomposed design, which allows these processes to be pre-computed and their results to be pre-stored.
$\bullet$ Addressing Limitations 3: The dynamic and evolving nature of business requirements can be naturally addressed in Step 3, similar to traditional DLRMs. Moreover, these dynamic requirements can be unified as various conditions, which aligns with our design for condition tokens. This means that regardless of changes in requirements, they can be consistently characterized as ¡condition token, item token¿, enabling continuous training of LUM to adapt to new requirements.
$\bullet$ Addressing Limitations 4: In an industrial setting, the DLRM backbone in Step 3 can be set as the online model. This setup allows for the easy utilization of pre-existing industrial knowledge. Furthermore, the model can continuously benefit from advancements in DLRM technology.
@@ 5.1. Experimental setting
Training Details. To ensure a fair comparison, we adhere to the following training configurations. By default, we maintain a similar configuration for transformer-style models, including LUM, HSTU, and SASRec, to ensure comparable model sizes. The configurations for other DLRMs are set according to the recommendations provided in their respective original papers. All models are trained from scratch using the same set of features. The sequence length is set to 256 for public datasets and 4096 for the industrial dataset. For LUM, the backbone of DLRMs in step 3 is configured as SIM for ranking tasks and EDB for retrieval tasks in public datasets. For the industrial dataset, the backbone is set as the online model currently deployed in production.
@@ 5.2. Performance on Recommendation tasks
5.2.1. Performance on public datasets
We first conduct experiments on public datasets to evaluate the performance of LUM. The performance metric reported is AUC [^1] Note 0.001 absolute AUC gain is regarded as significant for the ranking task(Zhou et al., 2018). The results are summarized in Table 2. From the results, we observe that LUM achieves significant improvements across all datasets. This indicates that the three-step paradigm based LUM effectively captures a wide range of user interests and enhances the predictive accuracy of DLRMs.
@@ 5.2.1. Performance on public datasets
5.2.2. Performance in industrial setting
In this section, we compare the performance of LUM against both DLRMs and E2E-GRs in an industrial setting. For the baseline, we use the online model in our applications, which follows an Embedding+MLP architecture for ranking and a two-tower architecture for retrieval. We report AUC for ranking and Recall@K (R@K) for retrieval. For LUM, the backbone DLRM in Step 3 is also set to the online model. Additionally, we include traditional state-of-the-art DLRMs (SASRec (Kang and McAuley, 2018) and DIN (Zhou et al., 2018)) and an E2E-GR (HSTU (Zhai et al., 2024)) as baselines for comparison. To ensure a fair comparison, all models use the same set of features. We set the maximum sequence length to 4096 for all models and train them from scratch. The results are summarized in Table 3. LUM achieves significant improvements over the best baseline, with a +0.0176 increase in AUC, a +0.0133 increase in R@10, and a +0.0134 increase in R@50. The substantial improvements can be attributed primarily to the generative-to-discriminative design of our proposed paradigm.
@@ 5.2.2. Performance in industrial setting
Table 5. Results of effectiveness evaluation.
Model Training Mode AUC
E2E-GR scratch 0.7334
DLRM scratch 0.7338
LUM scratch 0.7514
DLRM (feature) warming up 0.7541
DLRM (param) warming up 0.7525
DLRM (param+feature) warming up 0.7777
LUM (feature) warming up 0.7659
LUM (param) warming up 0.7620
LUM (param+feature) warming up 0.7794
LUM (w/o condition token) scratch 0.7416
LUM (multi-conditions) scratch 0.7545
LUM (direct feature) scratch 0.7402
LUM (direct feature+interest matching) scratch 0.7514
@@ 5.3. Effectiveness Evaluation
Impact of warming up setting in industrial applications. In industrial applications, online models typically incorporate complex feature engineering and are continually trained on billions of data per day. Existing E2E-GRs often ignore the knowledge derived from these online models, which can significantly challenge their performance, especially in applications where the online model has been developed over several years or even decades. LUM, however, offers excellent compatibility, allowing it to be trained in a warming up setting. In this section, we investigate the impact of this compatibility by implementing various versions of LUM, and comparing them with an E2E-GR (HSTU) and a DLRM (online model). We use AUC on an industrial dataset as the performance metric.
The results are shown in Table 5 (Note ”scratch” refers to models trained from scratch with the same set of features. ”feature” indicates the use of complex feature engineering. ”param” refers to initializing model parameters from the continually daily trained online model). From Table 5, we can draw the following conclusions: (1) Compared to LUM, the models under warming up settings, including LUM (feature), LUM (param), and LUM (feature+param), achieve noticeable gains in AUC, ranging from +0.0106 to +0.028. This highlights the significant benefit of leveraging existing knowledge from the online model. (2) While HSTU outperforms the DLRM in scratch setting, it still exhibits a large performance gap (0.7334 vs. 0.7777) compared to the DLRM (param+feature). This underscores the importance of being compatible with the online model. The limited compatibility of E2E-GRs may hinder their practical deployment in real-world applications.
@@
Table 6. The impact of group query and packing. Imp. refers to the relative improvements.
@@
Impact of packing and group query. Packing and group query are designed to accelerate the processes in Step 1 and Step 2, respectively. Table 6 demonstrates the efficiency gains of these strategies in an industrial setting, where Step 1 and Step 2 can be sped up by 82% and 78%, respectively.
@@ 5.4. Efficiency Evaluation
Training Efficiency. Figure 5 (a) illustrates the training time costs of different models on one day’s data in an industrial setting. For E2E-GR, we followed HSTU’s (Zhai et al., 2024) optimization and trained it from impression-level to user-level. For DLRM, we used the online model in our application as the baseline. For LUM, the backbone DLRM is also set as the online model. Since Steps 1 and 2 in LUM can be pre-processed (4.4), their time costs are excluded when training the downstream model in Step 3. The sequence length is set to 4096, and the model sizes of LUM and E2E-GR range from 0.5 billion to 14 billion parameters. All models are trained on 128 GPUs. Additionally, the lower bound time cost is set to 24 hours, as continuous training in practice requires models to process daily data within this timeframe. Overall, LUM achieves similar training time costs to DLRM and remains relatively insensitive to model size due to its decomposed three-step paradigm. This characteristic unlocks the potential for scaling laws during training in industrial settings. In contrast, E2E-GR is 12$\times$ to 98$\times$ slower than LUM. None of the E2E-GR models (across different sizes) meet the training cost requirement of completing within 24 hours. To match the throughput of LUM, E2E-GR would require 12$\times$ to 98$\times$ GPUs, and to meet the lower bound requirement, it would need 2$\times$ to 18$\times$ GPUs.
Serving Efficiency. Figures 5 (b) and (c) illustrate the latency of different models during online serving. For E2E-GR, we followed HSTU’s M-FALCON implementation (Zhai et al., 2024)  [^2] Note that, unlike LLMs where the trained parameters are seldom updated during serving, in industrial settings, the parameters of the ranking model are updated in real-time, meaning results or intermediate results cannot be shared across requests. Therefore, the cache strategy in M-FALCON is not applicable here. DLRM is set as the online model for both the baseline and the backbone ranking model of LUM. The latency of LUM during Step 3 is evaluated, as Step 1 and 2 can be pre-computed (Section 4.4). In Figure 5 (b), the sequence length is set to 4096, and the model sizes of LUM and E2E-GR range from 0.5 billion to 14 billion parameters. The latency constraint is set to less than 30 milliseconds (ms), and the number of ranking candidates is approximately 100 in our case. Overall, the latency of LUM is independent of the model size, allowing us to scale up LUM without violating the latency constraint. On the contrary, E2E-GR fails to provide timely responses (less than 30 ms) even for a small model (0.5 billion parameters). To further investigate, we attempted to reduce the sequence length of E2E-GR to meet the latency constraint (Figure 5 (c)). Disappointingly, the maximum sequence length that meets the latency constraint (¡30 ms) is only 64 when using a 14 billion parameter model, which is far from the typical setting (64 times smaller in our case). These results demonstrate that, while E2E-GR may exhibit scaling laws in offline performance, the practical application of large-scale models is severely limited in industrial settings due to latency constraints. In contrast, LUM’s decoupled architecture ensures consistent latency performance, making it a more feasible and scalable solution for real-time industrial applications.
@@ 5.5. Scaling Law for LUM
Following the protocols established in (Radford, 2018; Radford et al., 2019), we examine whether our LUM model complies to similar scaling law. Specifically, we focus on evaluating the scaling law with respect to model parameters and sequence length. For model parameters, to evaluate the impact of model parameters, we maintain a constant sequence length of 4096 and train models across a range of sizes, from 19 million to 7 billion parameters. For sequence length, to evaluate the impact of sequence length, we train models with varying sequence lengths, from 256 to 8192, under a fixed parameter size of 300 million. The results are plotted in Figure 6, where we observe a clear power-law scaling trend, consistent with previous findings (Zhai et al., 2024; Radford, 2018; Radford et al., 2019). The power-law scaling laws can be expressed as:
@@ 5.6. Online Results
To evaluate the effectiveness of LUM in an industrial setting, we implemented it in the sponsored search system of Taobao, the largest e-commerce platform in China. As discussed in Section 4.4, our evaluation process involves several key steps: (1) Offline, we first pretrained LUM and pre-triggered the necessary knowledge under various conditions. (2) The responses generated by LUM in Step 2 are stored, allowing them to be directly utilized during online serving. This pre-computation helps meet the strict latency constraints required in real-time industrial applications. (3) Finally, we conducted online A/B experiments to test LUM in the ranking task. Key performance metrics, CTR and RPM (Revenue Per Mile), demonstrated a significant improvement of 2.9% and 1.2% respectively. These findings highlight the practical benefits of LUM, demonstrating its ability to improve user engagement and business outcomes in large-scale e-commerce platforms.

### v2

@@ 4.3. Step 3: Knowledge Utilization in DLRMs
To enhance existing DLRMs, we propose two strategies (see Figure 3 (c)): (1) Direct Feature Incorporation: We integrate the outputs $o^{i}_{q_{n}}$ and the embedding $e^{i}_{i}$ of target item $i$ as fixed additional features into the DLRMs. This approach leverages the rich representations derived from the next-condition items directly. (2) Interest Matching via Similarity Measurement: We assess the alignment between the target item $i$ and user interests by computing the similarity $sim(o^{i}_{q_{n}},e^{i}_{i})$. This similarity score quantifies how well the target item matches the context provided by the next-condition items.
@@
4.4. Industrial Deployment
In this section, we describe the deployment of LUM in Taobao’s sponsored search system (Figure 5), where LUM is primarily integrated into the ranking stage. The core idea is based on the three-step paradigm: by decoupling the steps, we enable maximal parallelization across training and inference. During online serving, Step 2 (knowledge querying) is executed in real time during the early phase of user request processing—in parallel with retrieval and pre-ranking stage. This design decouples LUM inference from the strict latency constraints of the final ranking stage, allowing sufficient time for computation. We build a centralized cache system that serves two purposes: (1) it stores real-time inference results from knowledge querying; (2) it pre-loads offline-inferred user interests as fallbacks to ensure coverage under tight RT (response time) constraints. Then, the system retrieves the generated interests from the cache and feeds them to the DLRM in the ranking stage, while logging the results for offline training consumption. During offline training time, we adopt a synchronous training pipeline that aligns LUM pre-training with DLRM streaming training. For DLRM training, we directly consume logged user interests produced by online LUM inference—bypassing Step 1 and 2 entirely. This significantly improves training efficiency, a critical requirement for industrial-scale streaming systems. For LUM, we train it in parallel with DLRM. Once updated, the new LUM model is deployed online, and its inference outputs are logged to influence subsequent DLRM training—forming a closed-loop, co-evolutionary learning framework.
Figure 5. Online deployment.
@@ 4.5. Discussion
$\bullet$ Addressing Limitation 1: To address the first limitation, we design a generative-to-discriminative learning pipeline. In Step 1, the joint distribution $p(c_{1},i_{1},c_{2},i_{2},\dots,c_{L},i_{L})$ is learned via generative modeling over user behavior sequences. In Step 2, a next-condition-item prediction objective is introduced, enabling the model to activate relevant knowledge for downstream recommendation tasks. Finally, in Step 3, discriminative learning is performed to align with the requirements of discriminative CTR applications.
$\bullet$ Addressing Limitation 2: Efficiency demands arise from the need for downstream models to meet strict efficiency constraints. For E2E-GR, this implies tight latency and throughput requirements on the model itself. In contrast, our three-step paradigm naturally decouples pre-training and knowledge querying (Steps 1–2) from downstream deployment (Step 3). Thus, only Step 3 must satisfy real-time inference constraints. As illustrated in Figure 5, during offline training, LUM’s pre-training and knowledge querying run in parallel with DLRM’s streaming training—where DLRM consumes the querying outputs from logs—introducing no additional overhead to DLRM’s training pipeline. During online serving, knowledge is pre-computed and cached, leaving DLRM’s computation unaffected in ranking stage.
$\bullet$ Addressing Limitation 3: Like efficiency challenge, flexibility requirements are primarily imposed by downstream applications. For E2E-GR, both efficiency and flexibility must be jointly satisfied within a single model. In our framework, however, downstream flexibility needs are naturally handled by the DLRM in Step 3, highlighting the advantage of our decoupled design. Moreover, flexibility can also be supported during LUM pre-training: dynamic requirements (e.g., new user behaviors or business rules) can be unified as conditioning signals, consistent with our design of condition tokens. That is, any evolving requirement can be consistently represented as a tuple $\langle\text{condition token},\text{item token}\rangle$, enabling continuous adaptation of LUM through ongoing pre-training. Notably, unlike E2E-GR, LUM’s pre-training is not subject to strict efficiency constraints, allowing flexible integration of new conditions.
$\bullet$ Addressing Limitations 4: In our framework, the DLRM in Step 3 serves as the online ranking model. This design enables direct reuse of pre-existing industrial knowledge, including engineered features and learned representations. Moreover, the system can seamlessly incorporate ongoing advancements in DLRM architectures and training techniques, ensuring long-term maintainability and performance evolution.
@@ 5.1. Experimental setting
Training Configuration. To ensure a fair and rigorous comparison, we adopt a standardized training setup across all models. By default, transformer-based models—including LUM, HSTU, and SASRec—are configured to have comparable model sizes. Specifically, on public benchmark datasets, the number of parameters is set to approximately 30 million; on the industrial dataset, it is scaled to 0.3 billion to reflect real-world modeling demands. For other DLRM variants, model configurations follow the recommendations from their original papers. Notably, we try to increase the model size of these baselines but observe no further performance gains—consistent with findings reported in prior work (Zhou et al., 2018). By default, all models are trained from scratch using an identical feature set to ensure a controlled evaluation. The sequence length is set to 256 for public datasets and 4096 for the industrial dataset, reflecting differences in user behavior span and system requirements. The learning rate is set to 1e-3and 1e-4 for public and industrial datasets, respectively, and the batch size is fixed at 2048 for all experiments. Training is conducted using 32 GPUs for public datasets and 128 GPUs for the industrial dataset. For LUM, the DLRM backbone in Step 3 is instantiated as SIM for ranking tasks and EDB for retrieval tasks on public datasets. On the industrial dataset, the backbone is set to the online model currently deployed in production, ensuring alignment with real-world serving infrastructure.
@@ 5.2. Performance on Recommendation tasks
Performance on public datasets We first conduct experiments on public datasets to evaluate the performance of LUM. The performance metric reported is AUC [^2] Note 0.001 absolute AUC gain is regarded as significant for the ranking task(Zhou et al., 2018). The results are summarized in Table 2. From the results, we observe that LUM achieves significant improvements across all datasets. This indicates that the three-step paradigm based LUM effectively captures a wide range of user interests and enhances the predictive accuracy of DLRMs.
@@
Performance in industrial setting In this section, we compare the performance of LUM against both DLRMs and E2E-GRs in an industrial setting. For the baseline, we use the online model in our applications, which follows an Embedding+MLP architecture for ranking and a two-tower architecture for retrieval. We report AUC for ranking and Recall@K (R@K) for retrieval. For LUM, the backbone DLRM in Step 3 is also set to the online model. Additionally, we include traditional state-of-the-art DLRMs (SASRec (Kang and McAuley, 2018) and DIN (Zhou et al., 2018)) and an E2E-GR (HSTU (Zhai et al., 2024)) as baselines for comparison. To ensure a fair comparison, all models use the same set of features. We set the maximum sequence length to 4096 for all models and train them from scratch. The results are summarized in Table 3. LUM achieves significant improvements over the best baseline, with a +0.0176 increase in AUC, a +0.0133 increase in R@10, and a +0.0134 increase in R@50. The substantial improvements can be attributed primarily to the generative-to-discriminative design of our proposed paradigm.
@@ 5.3. Effectiveness Evaluation
Table 5. Results of effectiveness evaluation.
Model Training Mode AUC
E2E-GR training from scratch 0.7334
DLRM training from scratch 0.7338
LUM training from scratch 0.7514
DLRM (feature) incremental training 0.7541
DLRM (param) incremental training 0.7525
DLRM (param+feature) incremental training 0.7777
LUM (feature) incremental training 0.7659
LUM (param) incremental training 0.7620
LUM (param+feature) incremental training 0.7794
LUM (w/o condition token) training from scratch 0.7416
LUM (multi-conditions) training from scratch 0.7545
LUM (direct feature) training from scratch 0.7402
LUM (direct feature+interest matching) training from scratch 0.7514
Compatibility with Online Models In industrial applications, online models typically incorporate extensive feature engineering and are continuously updated using billions of interactions per day. Existing E2E-GRs often neglect the rich knowledge encoded in these production models—such as learned feature representations or parameter distributions—potentially undermining their performance, especially in mature systems where the online model has been refined over years or even decades. In contrast, LUM is designed with strong compatibility in mind, enabling seamless integration into incremental training pipelines. Specifically, LUM can inherit either (1) engineered features or (2) model parameters from the continually updated online model—allowing it to leverage historical knowledge and avoid relearning from scratch. In this section, we evaluate the impact of this compatibility by comparing various configurations of LUM against a representative E2E-GR (HSTU) and the production DLRM (i.e., the online model). We report AUC on an industrial dataset as the primary evaluation metric.
Results are summarized in Table 5 (“scratch”: trained from random initialization; “feature”: with production-grade features; “param”: initialized from the online model). We observe: (1) LUM variants under incremental training—LUM (feature), (param), and (feature+param) — outperform the scratch baseline by +0.0106 to +0.0280 in AUC, showing strong benefits from knowledge transfer. (2) While HSTU surpasses DLRM from scratch, it falls well short of the enhanced DLRM (feature+param) (0.7334 vs. 0.7777), underscoring the importance of integration with deployed systems. The limited adaptability of E2E-GRs to production knowledge may hinder real-world deployment.
@@
Table 6. The impact of group query and packing.
@@
Impact of packing and group query. Packing and group query are designed to accelerate Step 1 and Step 2, respectively. As shown in Table 6, these optimizations significantly reduce the computational cost of processing one day’s industrial data. Specifically, Step 1 and Step 2 achieve speedups of 82% and 78%, respectively, demonstrating their effectiveness in large-scale production environments.
@@ 5.4. Efficiency Evaluation
5.4.1. Training Efficiency
Setup. For E2E-GR, we follow HSTU (Zhai et al., 2024), training from impression-level to user-level. The DLRM baseline is our production model and also serves as the backbone for LUM in Step 3. Both LUM and E2E-GR use a sequence length of 4096 and scale from 0.5B to 14B parameters, trained on 128 GPUs. In industrial streaming training, models must process daily data within 24 hours. For E2E-GR (and DLRM), end-to-end training time defines the cost. For LUM, only Step 3 affects downstream latency due to the decoupled three-step paradigm (Section 4.5), so we report its cost as LUM-Step3. This setup is fair, realistic, and highlights the practical advantages of our framework.
[…]
5.4.2. Serving Efficiency
Setup: We follow the M-FALCON setup from HSTU (Zhai et al., 2024) for E2E-GR. [^3] Unlike LLMs with static parameters, industrial ranking models are updated in real time, making M-FALCON’s caching strategy inapplicable. The DLRM model serves as both the baseline and LUM backbone. We impose a strict inference latency budget of ¡30 ms per request, with  100 candidates ranked, reflecting real-world constraints. Input sequence length is 4096; model sizes for LUM and E2E-GR range from 0.5B to 14B. As discussed in Section 4.5, serving efficiency hinges on downstream latency. Since Steps 1 and 2 of LUM can be precomputed and cached, we evaluate only Step 3 latency—reported as LUM-Step3.
[…]
@@ 5.5. Scaling Law for LUM
Following the protocols established in (Radford, 2018; Radford et al., 2019), we examine whether LUM complies to similar scaling law. For sequence length, to evaluate the impact of sequence length, we train models with varying sequence lengths, from 256 to 8192, under a fixed parameter size of 300 million. The results are plotted in Figure 7, where we observe a clear power-law scaling trend, consistent with previous findings (Zhai et al., 2024; Radford, 2018; Radford et al., 2019). The power-law scaling laws can be expressed as:
@@ 5.6. Online Results
To evaluate the effectiveness of LUM in an industrial setting, we implemented it in the sponsored search system of Taobao, the largest e-commerce platform in China (see Section 4.4). We conducted online A/B experiments to test LUM in the ranking task. Key performance metrics, CTR and RPM (Revenue Per Mile), demonstrated a significant improvement of 2.9% and 1.2% respectively. These findings highlight the practical benefits of LUM, demonstrating its ability to improve user engagement and business outcomes in large-scale e-commerce platforms.